# Each to Their Own: Exploring the Optimal Embedding in RAG

### v2

@@ 4.2.1 Vanilla LLM and RAG
Table 1: Model Performance Comparison Across Different Embedding Methods: The four embedding models (1–4) correspond to all-MiniLM-L6-v2 Face , ModernBERT-large Warner et al. (2024), MathBERT Peng et al. (2021), and stsb-roberta-large Reimers and Gurevych (2019), respectively, which will remain consistent in the following tables. The last two columns represent the average accuracy of RAG using the four embedding models and the improvement compared to the vanilla LLM. These details will remain the same in the subsequent tables.
4.2.2 Results of Mixture-Embedding RAG method
Regarding the mixture-embedding RAG method (see Table 2), the average accuracies of the three LLMs did not perform well, compared to vanilla RAG. Llama-3.1-8B and OLMo-2-1124-7B demonstrated similar accuracy as vanilla RAG, with the differences ranging from -0.2% to 0.5%. Surprisingly, Qwen2.5-Math-7B showed a decrease of 5.5% compared to vanilla RAG.
@@ 4.2.2 Results of Mixture-Embedding RAG method
Table 2: Performance of Mixture-Embedding RAG: This table illustrates the performance when using between 2 and 4 different embedding models randomly (denoted as 2 Embs to 4 Embs), comparing the results with those of the vanilla LLM and vanilla RAG.
4.2.3 Results of Confident RAG method
While using different embedding models for RAG in multiple rounds, the results of Confident RAG method are shown in Table 3. Four main findings were obtained:
@@ 4.2.3 Results of Confident RAG method
Table 3: Accuracy Comparison Across Multi-RAG with Different Embedding Models: Avg($n$) denotes the average accuracy across different combinations of $n$ embedding models. Each line uses underline to indicate the best embedding combination within each LLM. Each row uses bold to signify the best metric for confidence evaluation. For performance comparison with Vanilla RAG, we use the average accuracy of all single embedding models as the baseline.
4.3 Analysis
4.3.1 Mixture-embedding RAG
@@
For general LLMs (e.g., Llama-3.1-8B and OLMo-2-1124-7B) without math fine-tuning, their internal math knowledge is limited, leading to lower accuracy in direct answer generation. In these cases, even noisy references retrieved by RAG are more reliable than the LLMs’ own outputs, as RAG at least provides partially correct information. However, while the mixture-embedding RAG method may optimize the retrieval ranking process and improve the quality of the references, the general LLMs’ capabilities prevent them from fully leveraging higher-quality references, resulting in performance similar to vanilla RAG. Additionally, if different embedding models return highly diverse references, directly combining the top-ranked documents may cause information overload or contextual confusion, negating the potential benefits of mixture-embedding method. Therefore, the performance of general LLMs matches that of vanilla RAG rather than surpassing it.
@@
On the other hand, for the LLMs that have been fine-tuned based on math corpora, vanilla RAG may result in smaller improvements, and lower-quality references can lead to poorer answer performance. For these types of LLMs, the mixture-embedding method may introduce additional noise in mathematical contexts, resulting in lower accuracy compared to vanilla RAG. The decline in accuracy may be caused by several factors: (1) Mathematical symbols and formulas may vary drastically across embedding models, making similarity calculations unstable. (2) Different models may encode math terms differently, causing the top-ranked reference to be suboptimal. (3) An embedding model might incorrectly rank an irrelevant math material highly, while better references from other models are ignored. If the LLM generates hallucinated answers based on incorrect references, its performance can degrade below that of the vanilla LLM. (4) Information overload or contextual confusion may also occur, similar to what happens with general LLMs.
4.3.2 Confident RAG
@@
As shown in Figure 2, there exists a positive correlation between confidence and accuracy. Therefore, the Confident RAG method improves overall accuracy by integrating multiple embedding models to generate answers and selecting the highest-confidence results using the most effective metric. This process effectively filters out low-confidence incorrect answers. Meanwhile, the combined effect of RAG and confidence filtering enhances the robustness, leading to significant improvements compared to vanilla LLMs. When employing the optimal confidence metric, all LLMs achieved an accuracy increase of nearly 10%, demonstrating the method’s universality. In the experiments, when the number of embedding models N > 3, the accuracy improvement became limited, likely due to redundant or noisy retrievals introduced by additional models. At N=3, the method achieved an optimal balance between diversity and computational efficiency. Increasing the number of models further yields only marginal benefits.
@@
5 Conclusion
In this paper, we proposed to enhance the response performance of LLMs with RAG utilizing multiple embedding models. We introduced two related approaches: Mixture-Embedding RAG and Confident RAG. Mixture-Embedding RAG sorts and selects retrieved references from different embedding models based on their standardized similarity before guiding LLM. However, its results showed no improvement over vanilla RAG, possibly due to factors such as information overload, contextual confusion and the capabilities of LLMs. In contrast, Confident RAG selects the highest-confidence response from multiple RAG outputs, each generated using a different embedding model. This approach demonstrated superior performance, with an average improvement of approximately 10% while utilizing the best confidence metric, compared to vanilla LLM. Self-certainty and distributional perplexity (DP) were identified as the most effective metrics for measuring model confidence. The number of embedding models (N) can be determined by researchers based on their specific conditions, as no evidence suggesting that a larger N necessarily yields a better performance. The consistent and stable results from Confident RAG indicate that it can serve as a plug-and-play method for enhancing LLM performance across various domains.

### v3

@@ IV-B Experiment Result
[…]
  • The four embedding models (1–4) correspond to all-MiniLM-L6-v2 [all-mini], ModernBERT-large [modernbert], MathBERT [peng2021mathbert], and stsb-roberta-large [reimers-2019-sentence-bert], respectively, which will remain consistent in the following tables. The last two columns represent the average accuracy of RAG using the four embedding models and the improvement compared to the vanilla LLM. These details will remain the same in the subsequent tables.
TABLE II: Sample generation results for an item
Question Correct Model Answer Confidence Accuracy
Answer (self-certainty)
Lee used to be able
to run the 400-meter
hurdles two seconds
faster than Gerald
would run the 400-
meter hurdles. But
Gerald changed his
diet, which improved
his speed by 10%.
If Lee runs the 400-
meter hurdles in 38
seconds, how fast
can Gerald, with his
improved diet, run
the 400-meter hurdles,
in seconds?
36 Vanilla LLM 34.2 / 0
Vanilla RAG
Emb1 34.2 9.1794 0
Emb2 36 16.7625 1
Emb3 36 15.9693 1
Emb4 34.2 9.3341 0
Mixture-embedding RAG
2 Embs 36 / 1
3 Embs 34.2 / 0
4 Embs 34.2 / 0
Confident RAG
1,2 36 16.7625 1
1,3 36 15.9693 1
1,4 34.2 9.3341 0
2,3 36 16.7625 1
2,4 36 16.7625 1
3,4 36 15.9693 1
1,2,3 36 16.7625 1
1,2,4 36 16.7625 1
1,3,4 36 15.9693 1
2,3,4 36 16.7625 1
1,2,3,4 36 16.7625 1
  • The LLM used in this example is Qwen2.5-Math-7B. The four embedding models (1-4) are the same as the previous table. The Mixture-embedding RAG results illustrate the performance when using retrieved information from 2 to 4 different embedding models randomly (denoted as 2 Embs to 4 Embs). The Confident RAG results illustrate the performance when using different combinations of multiple embedding models and using self-certainty as the confidence metric. Accuracy: 0=incorrect; 1=correct.
IV-B2 To what extent do Mixture-Embedding RAG enhance accuracy than vanilla LLM and RAG in mathematical problem-solving?
The Mixture-Embedding RAG method aimed to solve this instability by fusing retrieved documents. However, as shown in Table III, the average accuracies of the three LLMs did not perform well compared to vanilla RAG. Llama-3.1-8B and OLMo-2-1124-7B demonstrated similar accuracy as vanilla RAG, with the differences ranging from -0.2% to 0.5%. Surprisingly, Qwen2.5-Math-7B showed a decrease of 5.5% compared to vanilla RAG.
The example in Table II reveals the potential pedagogical danger of this approach. While using 2 embedding models (2 Embs) yielded the correct answer (36), fusing 3 or 4 embedding models (3 Embs, 4 Embs) resulted in incorrect answer (34.2). This suggests that the fusion process can sometimes amplify noisy or incorrect information, leading the LLM to a wrong conclusion and therefore resulting in a critical failure for educational tools designed to build correct conceptual understanding.
TABLE III: Performance of Mixture-Embedding RAG
LLM Model Mix-Embedding RAG
2 Embs 3 Embs 4 Embs Avg v.s. Vanilla LLM v.s. Vanilla RAG
[…]
  • This table illustrates the performance when using retrieved information from 2 to 4 different embedding models randomly (denoted as 2 Embs to 4 Embs), comparing the results with those of the vanilla LLM and vanilla RAG.
IV-B3 To what extent do Confident RAG enhance accuracy than vanilla LLM and RAG in mathematical problem-solving?
The Confident RAG method demonstrated superior and consistent performance, addressing the core instability of previous approaches. After using different embedding models for RAG in multiple rounds, the results of Confident RAG method are shown in Table IV. Four main findings were obtained:
@@ 2 Embs 36 / 1
The practical mechanism behind this success is clearly demonstrated in our running example (Table II). For the hurdle problem, Confident RAG generated multiple candidate answers. Crucially, the self-certainty metric correctly assigned a high score (16.76) to the correct answer (36) and significantly lower scores (9.3) to the incorrect one (34.2). In every combination of embedding models, we selected the answer with the high-confidence, which can largely minimizing the risk of propagating pedagogical errors and present the most reliable solution to the students.
@@
TABLE IV: Accuracy Comparison Across Multi-RAG with Different Embedding Models
LLM Emb. Model AvgLogP Self-certainty Gini Entropy DP
Qwen2.5-Math-7B [yang2024qwen2] 1,2 82.0% 85.0% 81.8% 82.8% 83.6%
1,3 79.4% 83.4% 79.2% 79.6% 81.0%
1,4 81.4% 84.6% 81.6% 81.8% 82.2%
2,3 81.8% 84.4% 81.4% 81.8% 82.4%
2,4 81.4% 83.6% 81.0% 81.4% 82.2%
3,4 79.2% 82.2% 79.8% 79.8% 80.6%
Avg (n=2) 80.9% 83.9% 80.8% 81.2% 82.0%
1,2,3 79.4% 85.0% 79.0% 79.8% 81.6%
1,2,4 79.6% 85.0% 79.6% 80.6% 82.0%
1,3,4 79.0% 84.8% 79.6% 80.0% 80.8%
2,3,4 79.2% 84.4% 79.0% 79.8% 81.0%
Avg (n=3) 79.3% 84.8% 79.3% 80.1% 81.4%
1,2,3,4 78.2% 84.8% 78.4% 79.2% 80.6%
Avg (n=2,3,4) 80.1% 84.3% 80.0% 80.6% 81.6%
@@
[…]
Llama-3.1-8B [grattafiori2024llama] 1,2 27.2% 27.2% 27.0% 27.0% 27.2%
1,3 26.8% 26.8% 26.6% 27.0% 27.2%
1,4 26.8% 26.4% 26.4% 26.4% 26.4%
2,3 23.2% 24.0% 24.0% 24.0% 23.6%
2,4 26.6% 26.8% 26.4% 26.0% 26.6%
3,4 25.8% 26.4% 26.6% 26.6% 26.4%
Avg (n=2) 26.1% 26.3% 26.2% 26.2% 26.2%
1,2,3 28.6% 29.2% 28.8% 29.4% 29.4%
1,2,4 28.8% 28.6% 28.0% 28.0% 28.2%
1,3,4 27.4% 27.4% 27.2% 27.6% 27.6%
2,3,4 25.8% 26.4% 26.0% 26.0% 26.4%
Avg (n=3) 27.7% 27.9% 27.5% 27.8% 27.9%
1,2,3,4 27.8% 27.6% 27.0% 27.4% 27.6%
Avg (n=2,3,4) 26.8% 27.0% 26.7% 26.9% 27.0%
[…]
OLMo-2-1124-7B [olmo20242] 1,2 31.0% 31.2% 30.4% 31.2% 32.2%
1,3 29.8% 29.6% 29.8% 30.0% 30.0%
1,4 29.4% 29.2% 28.6% 29.6% 31.0%
2,3 31.6% 30.8% 30.2% 31.4% 32.8%
2,4 32.6% 32.0% 31.0% 33.0% 33.8%
3,4 29.6% 30.2% 29.6% 30.6% 31.4%
Avg (n=2) 30.7% 30.5% 29.9% 31.0% 31.9%
1,2,3 32.6% 32.0% 31.4% 32.2% 33.2%
1,2,4 32.8% 32.8% 31.6% 33.2% 34.8%
1,3,4 31.2% 31.2% 30.6% 32.0% 34.6%
2,3,4 32.6% 32.2% 30.8% 33.6% 36.6%
Avg (n=3) 32.3% 32.1% 31.1% 32.8% 34.8%
1,2,3,4 32.8% 32.4% 31.0% 33.2% 35.8%
Avg (n=2,3,4) 31.5% 31.2% 30.5% 31.8% 33.3%
[…]
Average Avg (n=2) 45.9% 46.9% 45.6% 46.1% 46.7%
Avg (n=3) 46.4% 48.3% 46.0% 46.9% 48.0%
Avg (n=4) 46.3% 48.3% 45.5% 46.6% 48.0%
Avg (n=2,3,4) 46.1% 47.5% 45.7% 46.4% 47.3%
[…]
@@
  • Avg($n$) denotes the average accuracy across different combinations of $n$ embedding models. Each column uses underline to indicate the best embedding combination within each LLM. Each row uses bold to signify the best metric for confidence evaluation. For performance comparison with Vanilla RAG, we use the average accuracy of all single embedding models as the baseline.
V Discussion
V-A Mixture-embedding RAG Showed Limited Improvements from Vanilla RAG
@@
Our results indicate that Mixture-Embedding RAG did not provide the anticipated improvement. For general LLMs (e.g., Llama-3.1-8B and OLMo-2-1124-7B) without math fine-tuning, their internal math knowledge is limited, leading to lower accuracy in direct answer generation [lewkowycz2022solving]. In these cases, even noisy references retrieved by RAG are more reliable than the LLMs’ own outputs, as RAG at least provides partially correct information. However, while the mixture-embedding RAG method may optimize the retrieval ranking process and improve the quality of the references, the general LLMs’ capabilities prevent them from fully leveraging higher-quality references, resulting in performance similar to vanilla RAG. Additionally, if different embedding models return highly diverse references, directly combining the top-ranked documents may cause information overload or contextual confusion, negating the potential benefits of mixture-embedding method. Therefore, the performance of general LLMs matches that of vanilla RAG rather than surpassing it.
@@
On the other hand, for the LLMs that have been fine-tuned based on math corpora (e.g., Qwen2.5-Math-7B), vanilla RAG may result in smaller improvements, and lower-quality references can lead to poorer answer performance [lewis2020retrieval]. For these types of LLMs, the mixture-embedding method may introduce additional noise in mathematical contexts, resulting in lower accuracy compared to vanilla RAG. The decmidrule in accuracy may be caused by several factors: (1) Mathematical symbols and formulas may vary drastically across embedding models, making similarity calculations unstable. (2) Different models may encode math terms differently, causing the top-ranked reference to be suboptimal. (3) An embedding model might incorrectly rank an irrelevant math material highly, while better references from other models are ignored. If the LLM generates hallucinated answers based on incorrect references, its performance can degrade below that of the vanilla LLM. (4) Information overload or contextual confusion may also occur, similar to what happens with general LLMs.
V-B Confident RAG Demonstrated Consistent Improvements through Confidence Selection
@@
As shown in Figure 2, there exists a positive correlation between confidence and accuracy, which also aligns with the finding from [chen2025first]. Therefore, the Confident RAG method improves overall accuracy by integrating multiple embedding models to generate answers and selecting the highest-confidence results using the most effective metric. This process effectively filters out low-confidence incorrect answers. Meanwhile, the combined effect of RAG and confidence filtering enhances the robustness, leading to significant improvements compared to vanilla LLMs. When employing the optimal confidence metric, all LLMs achieved an accuracy increase of nearly 10%, demonstrating the method’s universality. In the experiments, when the number of embedding models N ¿ 3, the accuracy improvement became limited, likely due to redundant or noisy retrievals introduced by additional models. At N=3, the method achieved an optimal balance between diversity and computational efficiency. Increasing the number of models further yields only marginal benefits.
@@
V-C Practical Implications
The findings of this study lead to the following three practical implications.
First, our work on Confident RAG establishes a critical foundation for the future deployment of fully autonomous Agentic RAG systems in educational settings. Leverages multiple embedding models and a confidence-based selection mechanism, our results indicate that it can achieve an accuracy improvement of approximately 10% compared to the vanilla LLM when using a self-certainty or DP confidence metric.
For autonomous planning [huang2024understanding], the agentic system could dynamically determine the optimal number of embedding models to employ based on the confidence scores, and select specific embedding models from a broader bank according to their historical performance. For iterative refinement [huang2023towards], the system could automatically initiate additional rounds of retrieval using different embedding models if the initial results do not meet a predefined confidence standard, ensuring high-quality output. For long-term memory [zhang2025survey], the system could keep a learning record, remembering which embedding models work best for different kinds of math problems, allowing it to become smarter and more efficient over time.
Second, this research gives educators and schools a simple way to judge AI teaching tools. They should prioritize tools that demonstrate a methodology for cross-verification and confidence estimation. Our Confident RAG method is an example of this safer and more reliable approach, which is crucial for students’ learning.
Third, the Confident RAG method supports students’ independent learning and saves teachers’ time. With a more trustworthy AI, students can do their homework and explore topics on their own with less risk of learning from a wrong answer. This builds their confidence in self-directed learning. At the same time, it reduces the heavy burden on teachers. They will spend less time double-checking every AI-generated answer for mistakes, freeing them up to focus on actual teaching and helping students who need it most.
VI Limitations and Future Study
This study has two main limitations. First, the experiment was only designed for mathematics domain using the GSM8K benchmark and a corpus of mathematical textbooks. This design prevents generalization of our findings to other subjects in educational context. Future study should validate the Confident RAG framework across diverse academic domains. This will test its robustness and determine whether it is a truly general-purpose solution for educational AI.
Second, this study provides a validated methodological framework (Confident RAG) that researchers and developers can reference to build more reliable AI math assistants. However, we have not yet embedded this method into a complete, user-facing educational system. Therefore, its usability and practical impact in a real learning environment remain unexplored. Future study should build a functional AI math tutoring system that incorporates the Confident RAG framework. This will allow us to validate the method’s effectiveness not just on benchmark accuracy, but on its ability to support student learning through interactive use, and to refine its design based on user feedback from actual classrooms.
VII Conclusion
In this paper, we proposed two novel RAG-enhanced approaches using multiple embedding models to improve the mathematics question response performance of LLMs: (1) Mixture-Embedding RAG and (2) Confident RAG. Our findings clearly indicate that Confident RAG is a superior and robust solution, with an average improvement of approximately 10% while utilizing the best confidence metric, compared to vanilla LLM. It selects the highest-confidence response from multiple RAG outputs, each generated using a different embedding model. The consistent success of confidence metrics like self-certainty and DP provides a reliable mechanism for identifying correct answers. Crucially, the number of embedding models (N) can be determined by researchers based on their specific conditions, as no evidence suggesting that a larger N necessarily yields a better performance. In contrast, Mixture-Embedding RAG showed no improvement over vanilla RAG, by sorting and selecting retrieved references from different embedding models based on their standardized similarity before guiding LLM. The findings indicate that Confident RAG can be served as a plug-and-play method for enhancing LLM performance in mathematics problem-solving and building more accurate and reliable Agentic RAG system in mathematics education.